OpenStudy (anonymous):

Please help, The last question on my test! Will give medal and fan! The histogram shows the temperature, in degrees Celsius, of a city (x) on different number of days (y). Based on the histogram, which statement is correct about the shape of the histogram?

OpenStudy (anonymous):

It is a bell curve with a longer tail on the right because of the presence of more number of days with higher temperatures than with lower temperatures. It is a bell curve with a longer tail on the right because of the presence of more number of days with lower temperatures than with higher temperatures. It is a bell curve with a longer tail on the left because of the presence of more number of days with lower temperatures than with higher temperatures. It is a bell curve with a longer tail on the left because of the presence of more number of days with higher temperatures than with lower temperatures. ~The graph is attached~

OpenStudy (anonymous):

@Nurali @zepdrix can you help?

OpenStudy (anonymous):

@kropot72 @tHe_FiZiCx99

OpenStudy (anonymous):

@Orion1213 sorry to keep bothering you but can you help?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

It was wrong...

OpenStudy (anonymous):

i think it's the first...

OpenStudy (kropot72):

Total number of days with temperatures below zero = 4 + 3 + 2 + 1 = 10 Total number of days with temperatures above zero = 8 + 6 + 5 + 4 = 23

OpenStudy (anonymous):

The first one was wrong tho..

OpenStudy (kropot72):

@_Lissie_ Didn't you find the second choice was wrong?

OpenStudy (anonymous):

Yes, the second one would be too

OpenStudy (kropot72):

Then it must be the last one.

OpenStudy (anonymous):

got your point... you have a long tail in the left side because you got a sudden reduction in temperature considering T<5°...

OpenStudy (kropot72):

That's it!

OpenStudy (anonymous):

Thank you @kropot72

OpenStudy (kropot72):

You're welcome :)